Depth of bore
edge condition Above target Below target
coarse 15 10
moderate 21 24
smooth 49 81

The data from 200 machined parts are summarized as follows:

(a) What is the probability that a part selected has a moderate edge condition and a below-target bore depth?
Round your answer to two decimal places (e.g. 98.76).

P=

(b) What is the probability that a part selected has a moderate edge condition or a below-target bore depth?
Round your answer to two decimal places (e.g. 98.76).

P=

(c) What is the probability that a part selected does not have a moderate edge condition or does not have a below-target bore depth?
Round your answer to two decimal places (e.g. 98.76).

P=
Answer #1
Depth of bore
edge condition Above target Below target total
coarse 15 10 25
moderate 21 24 45
smooth 49 81 130
total 85 115 200

a)

P(moderate edge and below target bore  condition) = 24/200=0.12

b)

probability ( a part selected has a moderate edge condition or a below-target bore depth) = [ n(moderate)+n(below) - n(moderate and below) ] / n(total) = (45+115-24)/200=0.68

c)

probability ( a part selected does not have a moderate edge condition or does not have a below-target bore depth)=

=1 - P(a part selected has a moderate edge condition or a below-target bore depth) = 1-0.68 = 0.32